# GLEAM: Learning to Match and Explain in Cross-View Geo-Localization

### v2

@@ 3.3 Training Recipe
We propose a two-phase training scheme to effectively integrate knowledge across multiple modalities and views (Sec. 3.3.1). Given the substantial data requirements of foundational model training and the limited efficiency of the original Sample4Geo codebase, we implement targeted optimizations to significantly enhance training efficiency (Sec. 3.3.2).
@@ 3.3.1 Two-Phase Training Strategy
We note that a naive concatenation of multiple datasets for training from the outset poses significant challenges. Due to the backbone model’s limited matching capacity and the varying dataset sizes and difficulty levels, direct multi-dataset training often results in highly imbalanced performance. To mitigate this issue, we first train the model on a single dataset that is relatively large and of moderate difficulty, allowing the backbone to acquire fundamental CVGL capabilities before introducing more complex and diverse data. In our design, we first train the model on the VIGOR (Zhu et al., 2021) dataset for 40 epochs, followed by training on the concatenated datasets for an additional 40 epochs.
@@ 3.3.2 Training Efficiency Improvement
Similar to CLIP training, contrastive learning models typically require large batch sizes. The original Sample4Geo adopts PyTorch’s Data Parallel (DP) strategy, where GPU 0 distributes data, aggregates outputs, and performs gradient updates for the entire cluster. This centralized design creates a significant communication bottleneck on GPU 0, thus slowing the training process. Distributed Data Parallel (DDP) alleviates this issue by allowing each GPU to compute gradients locally and synchronizing them via an efficient all-reduce. However, contrastive learning relies on in-batch negatives, and small per-GPU batch sizes limit the number of negatives available per device, potentially degrading performance. This underscores the importance of maintaining sufficiently large per-GPU batch sizes even when using DDP. To tackle this issue, we adopt the strategy implemented in OpenCLIP (Ilharco et al., 2021). Before computing the contrastive loss, each GPU (rank) gathers the feature representations from all other GPUs. These features are then re-concatenated, effectively creating a unified global batch for the loss calculation on each device. This process preserves the gradient flow for proper backpropagation while providing every GPU with a much larger and more effective set of negative samples. For clarity, we provide pseudocode for the loss computation in Alg. 1 in Sec. A.1. This implementation achieves over 5-fold faster training speed while maintaining the original model accuracy.
@@ 4 GLEAM-X: A Multi-Image MLLM Explanation Benchmark
GLEAM-X extends GLEAM-C to a multi-image reasoning setting, where the model not only predicts matches between query and reference images but also generates interpretable explanations. In practice, GLEAM-C and GLEAM-X operate as an integrated pipeline: GLEAM-C first retrieves the best-matching reference image for a given query, and GLEAM-X then verifies whether this reference image truly matches the query while providing human-interpretable explanations. In this section, we present the data composition (Sec. 4.1), annotation procedure (Sec. 4.2), and evaluation protocol (Sec. 4.3) of the GLEAM-X benchmark.
@@ 4.1 Data Composition
Our training and testing data pairs are derived from the GLEAM-C training and test sets. To maintain balance across modalities and views, we perform data sampling to ensure that the amount of training data for each modality is roughly equal, as shown in Tab. 1. In the evaluation of GLEAM-X, we also maintain this balance across modalities. The selected query images for testing include 128 for MAP, 126 for SetVL-480K, 127 for University-1652, and 123 for VIGOR. For each query image, we select one positive reference image and one negative reference image. Using the corresponding labels, we generate explanation annotations in both English and Chinese with commercial MLLMs, specifically GPT-4o (Hurst et al., 2024) and Doubao-1.5-Thinking-Vision-Pro (Volcengine, 2025).
@@ 4.2 Data Annotation
To ensure the accuracy of the test set, we manually revise the 2k test pairs. We engage 4 human experts in Remote Sensing, each holding at least a master’s degree, to perform data annotation and revision. We conduct two annotation rounds: in each round, we first verify the correctness of the explanations generated by the commercial MLLMs and correct them if necessary; if the four experts cannot determine whether a query and reference image match, we replace the pair with a suitable one from the remaining test data and perform model-assisted generation followed by manual correction. This ensures that all test pairs are valid and unambiguous, providing a reliable basis for evaluation.
4.3 Evaluation Protocol
We evaluate the MLLM outputs on the test set from two perspectives. First, we extract the predicted match labels using regular-expression parsing and compute the matching accuracy. Second, we assess the quality of the generated explanations by measuring their semantic similarity to the annotated references with Sentence-BERT (Reimers & Gurevych, 2019). This dual evaluation framework ensures a comprehensive assessment, as it captures both the correctness of structured match predictions and the semantic fidelity of explanatory outputs.
5 Experiments
In this section, we design a series of experiments to evaluate the unified GLEAM-C model (Sec. 5.1) and the GLEAM-X benchmark (Sec. 5.2). All training experiments are conducted on a cluster equipped with 10 NVIDIA RTX 4090 D GPUs, each with 48 GB of memory.
5.1 Evaluation results of GLEAM-C
For GLEAM-C, we compare the unified model with single-dataset training and evaluate the data integration strategy (Sec.5.1.1). We further assess improvements in training efficiency (Sec.5.1.2) relative to prior SOTA training methods.
@@ 5.1 Evaluation results of GLEAM-C
ConvNeXt-B-384
Single-dataset
Sample4Geo
w/o GPS
VIGOR 40 76.60 95.07 96.83 99.64 88.29
@@
From-scratch
VIGOR 40 73.37 93.43 95.66 99.61 84.43
@@
Two-phase
(VIGOR $\rightarrow$ Merge)
VIGOR 40 75.66 94.51 96.43 99.65 86.80
@@
From-scratch
VIGOR 40 69.44 92.81 95.43 99.63 83.18
@@
Two-phase
(VIGOR $\rightarrow$ Merge)
VIGOR 40 75.96 95.46 97.17 99.68 89.44
@@
Table 2: Image retrieval accuracy across backbones and training strategies. For University-1652, the last evaluation metric is AP; for all other datasets, it is Hit Rate. The first three rows (highlighted in gray) are directly taken from the corresponding papers with GPS (if applicable, marked with ) and DSS sampling for reference. It can be observed that using a two-phase training strategy yields relatively stronger performance. The unified GLEAM-C model achieves results comparable to or even exceeding those obtained by single-dataset training (Sample4Geo w/o GPS-Sampling).
@@
[…]
5.1.1 From Single Dataset to Multi-Dataset Integration
As in most CVGL studies, we first train on single-view and single-modality datasets. Training is conducted separately on the VIGOR (same), University-1652, MAP, and SetVL-480K (N = 1) datasets. Model performance is evaluated using image retrieval accuracy, with University-1652 following the Drone2Sat protocol. We adopt ConvNeXt-B-384 (88.6M) and PE-Core-L14-336 (0.32B) as the image encoders. The results for single-dataset training are presented in Tab. 2. Since some datasets lack GPS metadata, we do not use GPS sampling in all our model training, which may result in a slight decrease in accuracy. For reference, we also report the original metrics from the respective papers, with GPS (if applicable) and DSS sampling (Deuser et al., 2023b; Wu et al., 2024a).
2) Compared with single-dataset training, our two-phase training strategy yields the foundational GLEAM-C model, which attains comparable or superior performance across all evaluated datasets.
5.1.2 Training Efficiency Evaluation
We evaluate training efficiency on the VIGOR dataset by comparing DP and DDP strategies. Following the experiment settings of Sample4Geo (Deuser et al., 2023b), we adopt the ConvNeXt-B-384 backbone and use the original Sample4Geo code for DP training (without GPS-based sampling). As shown in Tab. 3, both approaches achieve comparable image retrieval accuracy across all Recall@k and Hit Rate metrics. However, DDP demonstrates a substantial (5-fold) speed advantage, reducing the time per training epoch from 521 seconds to 92 seconds, highlighting its effectiveness for accelerating large-scale CVGL model training.
5.2 Evaluation results on GLEAM-X
@@
We evaluate both commercial models and open-source models on the GLEAM-X benchmark. For commercial models, we regenerate answers on the test set using GPT-4o and Doubao-1.5-Thinking-Vision-Pro without providing ground-truth labels. For open-source models, we compare three variants of Qwen2.5-VL-3B-Instruct: the original model, the version fine-tuned with LLM-explanation supervision, and a version trained solely with label-only supervision. Specifically, for the explanation-based variant, we follow the procedure in Sec. 4.1, concatenating labels and explanations during training. We train separately on the 200k explanations generated by GPT-4o and Doubao-1.5 for 1 epoch. Conversely, the label-only variant uses only labels (e.g., [[1]]) for training.
After inference on the test set, we decouple matching predictions and explanations through regularization to isolate their respective contributions. To provide a comprehensive evaluation of GLEAM-X, we perform detailed analyses from three perspectives: matching accuracy (Sec.5.2.1), semantic accuracy (Sec.5.2.2), and positive/negative sample analysis (Sec. 5.2.3).
@@ 5.2.1 Matching Accuracy
The matching accuracy results are presented in Tab. 4. Commercial models achieve a relatively strong performance of around 80%, with GPT-4o outperforming Doubao-1.5 on average. In contrast, the original Qwen2.5-VL-3B-Instruct model exhibits very low accuracy, particularly in the Chinese setting, where its accuracy is only 50.50%. After training, all model variants obtain accuracy surpassing that of the commercial MLLMs. Notably, the label-only supervision model achieves the highest accuracy, likely because its training targets are extremely short (e.g., [[0]] or [[1]]), which reduces the learning difficulty for the model.
@@ 5.2.2 Semantic Analysis of Explanations
Tab. 5 reveals the similarity distributional differences across models. Commercial MLLMs concentrate heavily in the high-similarity range (0.8–1.0) for English, but shift toward 0.6–0.8 in Chinese. The original Qwen2.5-VL-3B-Instruct places a substantial proportion of predictions in the mid-similarity bins (0.6–0.8) for both languages, indicating weaker semantic alignment with ground-truth explanations. Fine-tuning with LLM-explanation supervision shifts the distribution upward, with both +GPT-4o and +Doubao-1.5 variants peaking in the 0.8–1.0 range. By contrast, the label-only supervision model collapses almost entirely into the lowest bin (0.0–0.2), confirming its inability to generate semantically meaningful explanations despite improvements in matching accuracy.
5.2.3 Analysis of Positive and Negative Pair Samples
Model Language Pos Acc Neg Acc Pos-Neg Pos Sim Neg Sim Pos-Neg
GPT-4o 70.83 91.67 -20.84 0.8296 0.8401 -0.0105
Doubao-1.5 66.87 88.10 -21.23 0.8100 0.8241 -0.0141
Qwen2.5-VL-3B-Instruct 23.81 90.87 -67.06 0.7616 0.7675 -0.0059
   +label only 93.85 91.67 2.18 0.0851 0.1111 -0.0260
   +GPT-4o 93.25 87.50 5.75 0.8399 0.8411 -0.0012
   +Doubao-1.5 EN 90.08 89.48 0.60 0.8048 0.8195 -0.0147
Model Language Pos Acc Neg Acc Pos-Neg Pos Sim Neg Sim Pos-Neg
GPT-4o 73.02 94.25 -21.23 0.7337 0.7468 -0.0131
Doubao-1.5 76.19 82.14 -5.95 0.7435 0.7555 -0.0120
Qwen2.5-VL-3B-Instruct 1.19 99.80 -98.61 0.7265 0.7533 -0.0268
   +label only 93.85 91.67 2.18 0.1471 0.1481 -0.0010
   +GPT-4o 90.87 89.48 1.39 0.8211 0.8267 -0.0056
   +Doubao-1.5 ZH 84.52 91.07 -6.55 0.7858 0.7989 -0.0131
Table 6: Positive and negative sample accuracy (Pos/Neg Acc) and similarity (Pos/Neg Sim) for different models on GLEAM-X in English and Chinese. The Pos-Neg columns report the difference between positive and negative samples. Vanilla Qwen2.5-VL-3B-Instruct shows a strong bias toward predicting non-matching pairs, while LLM-explanation supervision (+GPT-4o, +Doubao-1.5) mitigates this bias and improves alignment with ground-truth explanations.
To further analyze model behavior on positive and negative samples, we examine both matching accuracy and semantic similarity in Tab. 6. Vanilla MLLMs, including both commercial models and Qwen2.5-VL-3B-Instruct, exhibit a bias toward negative samples. In particular, Qwen2.5-VL-3B-Instruct achieves very low positive accuracy while maintaining high negative accuracy, indicating that the original model tends to predict image pairs as non-matching. Fine-tuning with LLM-explanation supervision (+GPT-4o, +Doubao-1.5) substantially mitigates this bias, raising positive accuracy to over 90% in English and over 84% in Chinese, while still keeping negative accuracy high. Correspondingly, positive and negative similarity are more balanced. Overall, LLM-explanation supervision effectively improves both prediction balance and explanation quality across languages. In contrast, the label-only supervision variant achieves high positive accuracy but fails to generate semantically meaningful explanations, as evidenced by extremely low similarity scores.

### v3

@@ 3.3 Training Recipe
We propose a novel two-phase training scheme to integrate knowledge across multiple modalities and views (Sec. 3.3.1). Given the substantial data requirements of foundational model training and the limited efficiency of the original Sample4Geo codebase, we implement targeted optimizations to significantly enhance training efficiency (Sec. 3.3.2).
@@ 3.3.1 Two-Phase Training Strategy
We note that a naive concatenation of multiple datasets for training poses significant challenges. Due to the backbone model’s limited matching capacity and the varying dataset sizes and difficulty levels, direct multi-dataset training often results in highly imbalanced performance. To mitigate this issue, we first novelly train the model on a single dataset that is relatively large and of moderate difficulty, allowing the backbone to acquire fundamental CVGL capabilities before introducing more complex and diverse data. In our design, we first train the model on the VIGOR (Zhu et al., 2021) dataset for 40 epochs, followed by training on the concatenated datasets for an additional 40 epochs.
@@ 3.3.2 Training Efficiency Improvement
Similar to CLIP training, contrastive learning models typically require large batch sizes. The original Sample4Geo adopts PyTorch’s Data Parallel (DP) strategy, where GPU 0 distributes data, aggregates outputs, and performs gradient updates for the entire cluster. This centralized design creates a significant communication bottleneck on GPU 0, thus slowing the training process. Distributed Data Parallel (DDP) alleviates this issue by allowing each GPU to compute gradients locally and synchronizing them via an efficient all-reduce. However, contrastive learning relies on in-batch negatives, and small per-GPU batch sizes limit the number of negatives available per device, potentially degrading performance. This underscores the importance of maintaining sufficiently large per-GPU batch sizes even when using DDP. To tackle this issue, we adopt the strategy implemented in OpenCLIP (Ilharco et al., 2021). Before computing the contrastive loss, each GPU (rank) gathers the feature representations from all other GPUs. These features are then re-concatenated, effectively creating a unified global batch for the loss calculation on each device. This process preserves the gradient flow for proper backpropagation while providing every GPU with a much larger and more effective set of negative samples. We provide pseudocode for the loss computation in Alg. 1 in Sec. A.1. This implementation achieves over 5-fold faster training speed while maintaining the original model accuracy.
@@ 4 GLEAM-X: A Multi-Image MLLM Explanation Benchmark
GLEAM-X extends GLEAM-C to a multi-image reasoning setting, where the model not only predicts matches between query and reference images but also generates interpretable explanations. In this section, we present the data composition (Sec. 4.1), annotation procedure (Sec. 4.2), evaluation protocol (Sec. 4.3), and how it can be combined with GLEAM-C as an integrated pipeline (Sec. 4.4).
@@ 4.1 Data Composition
Our training and testing data pairs are derived from the GLEAM-C training and test sets. To maintain balance across modalities and views, we perform data sampling to ensure that the amount of training data for each modality is roughly equal, as shown in Tab. 1. In the evaluation of GLEAM-X, we also maintain this balance across modalities. The selected query images for testing include 128 for MAP, 126 for SetVL-480K, 127 for University-1652, and 123 for VIGOR. For each query image, we select one positive and one negative reference image. Using the corresponding labels, we generate explanation annotations in both English and Chinese with commercial MLLMs, specifically GPT-4o (Hurst et al., 2024) and Doubao-1.5-Thinking-Vision-Pro (Volcengine, 2025). Notably, to ensure linguistic consistency and semantic equivalence in the bilingual benchmark, we generate both language versions simultaneously in a single prompt, explicitly requiring the model to provide corresponding explanations in English and Chinese.
@@ 4.2 Data Annotation
To ensure the accuracy of the test set, we manually revise the 2k test pairs. We engage 4 human experts in Remote Sensing, each holding at least a master’s degree, to perform data annotation and revision. We conduct two annotation rounds: in each round, we first verify the correctness of the explanations generated by the commercial MLLMs and correct them if necessary; if the four experts cannot determine whether a query and reference image match, we replace the pair with a suitable one from the remaining test data and perform model-assisted generation followed by manual correction. This ensures that all test pairs are valid and unambiguous, providing a reliable basis for evaluation. More details of human annotation are provided in Sec. A.2.
Table 2: Image retrieval accuracy across backbones and training strategies. For University-1652, the last evaluation metric is AP; for all other datasets, it is Hit Rate. The first three rows (highlighted in gray) are directly taken from the corresponding papers with GPS (if applicable, marked with ) and DSS sampling for reference. Using a two-phase training strategy yields relatively stronger performance. The unified GLEAM-C model achieves results comparable to or even exceeding those obtained by single-dataset training.
@@
ConvNeXt-B-384 Single-dataset Sample4Geo w/o GPS VIGOR 40 76.60 95.07 96.83 99.64 88.29
@@
From-scratch (Merge) VIGOR 40 73.37 93.43 95.66 99.61 84.43
@@
Two-phase (VIGOR $\rightarrow$ Merge) VIGOR 40 75.66 94.51 96.43 99.65 86.80
@@
From-scratch (Merge) VIGOR 40 69.44 92.81 95.43 99.63 83.18
@@
Two-phase (VIGOR $\rightarrow$ Merge) VIGOR 40 75.96 95.46 97.17 99.68 89.44
@@
[…]
@@
4.3 Evaluation Protocol
We evaluate the MLLM outputs on the test set from two perspectives. First, we extract the predicted match labels using regular-expression parsing and compute the matching accuracy. Second, we assess the quality of the generated explanations by measuring their semantic similarity to the annotated references with Sentence-BERT (Reimers and Gurevych, 2019). This dual evaluation framework ensures a comprehensive assessment, as it captures both the correctness of structured match predictions and the semantic fidelity of explanatory outputs.
4.4 GLEAM-C and GLEAM-X as a Pipeline
In real-world deployment, GLEAM-C and GLEAM-X operate as a two-stage pipeline that combines efficient image retrieval with interpretable verification, as illustrated in Fig. 1. Given a query image and a reference database, GLEAM-C first retrieves the most similar reference image through CVGL. GLEAM-X then verifies whether the query and retrieved reference correspond to the same geographic location, while providing human-interpretable explanations that identify key visual correspondences or discrepancies between the image pair. This integrated approach not only enhances retrieval reliability but also offers transparency through explainable reasoning, which is critical for safety-critical real-world applications such as autonomous navigation, search and rescue operations, and emergency response.
5 Experiments
In this section, we design a series of experiments to evaluate the unified GLEAM-C model (Sec. 5.1) and the GLEAM-X benchmark (Sec. 5.2). All training experiments are conducted on a cluster with 10 NVIDIA RTX 4090 D GPUs, each with 48 GB of memory. In addition, to assess practical deployment feasibility, we evaluate the models’ inference performance on an NVIDIA Jetson AGX Xavier for real-world edge computing scenarios in the Appendix (Sec. A.3).
5.1 Evaluation Results of GLEAM-C
For GLEAM-C, we compare the unified model with single-dataset training and evaluate the data integration strategy (Sec.5.1.1). We further assess improvements in training efficiency (Sec.5.1.2) relative to prior SOTA training methods. Additionally, we investigate cross-domain generalization performance and the rationale for selecting VIGOR as the initial training dataset (Sec. 5.1.3).
@@ 5.1 Evaluation Results of GLEAM-C
5.1.1 From Single Dataset to Multi-Dataset
As in most CVGL studies, we first train on single-view and single-modality datasets. Training is conducted separately on the VIGOR (same), University-1652, MAP, and SetVL-480K (N = 1) datasets. Model performance is evaluated using image retrieval accuracy, with University-1652 following the Drone2Sat protocol. We adopt ConvNeXt-B-384 (88.6M) and PE-Core-L14-336 (0.32B) as the image encoders. The results for single-dataset training are presented in Tab. 2. Since some datasets lack GPS metadata, we do not use GPS sampling in all our model training, which may result in a slight decrease in accuracy. For reference, we also report the original metrics from the respective papers, with GPS (if applicable) and DSS sampling (Deuser et al., 2023a; Wu et al., 2024b).
2) Compared with single-dataset training, our novel two-phase training strategy yields the foundational GLEAM-C model, which attains comparable or superior performance across all evaluated datasets.
5.1.2 Training Efficiency Evaluation
We evaluate training efficiency on the VIGOR dataset by comparing DP and DDP strategies. Following the experiment settings of Sample4Geo (Deuser et al., 2023a), we adopt the ConvNeXt-B-384 backbone and use the original Sample4Geo code for DP training (without GPS-based sampling). As shown in Tab. 3, both approaches achieve comparable image retrieval accuracy across all Recall@k and Hit Rate metrics. However, DDP demonstrates a substantial (5-fold) speed advantage, reducing the time per training epoch from 521 seconds to 92 seconds, highlighting its effectiveness for accelerating large-scale CVGL model training.
5.1.3 Discussion: Cross-Domain Generalization & Initial Dataset Selection
Beyond the above analysis, we further investigate two questions regarding our unified model training scheme: 1) What is the cross-domain generalization performance when trained on one modality and tested on unseen modalities? 2) Why VIGOR is chosen as the initial dataset in our two-phase training strategy? Detailed experimental analysis and results are provided in Sec. A.4 and Sec. A.5, respectively.
5.2 Evaluation Results on GLEAM-X
We evaluate both commercial models and open-source models on GLEAM-X. For commercial models, we regenerate answers on the test set using GPT-4o and Doubao-1.5-Thinking-Vision-Pro without providing ground-truth labels. For open-source models, we compare three variants of Qwen2.5-VL-3B-Instruct: the original model, the version fine-tuned with LLM-explanation supervision, and a version trained solely with label-only supervision. Specifically, for the explanation-based variant, we follow the procedure in Sec. 4.1, concatenating labels and explanations during training. We train separately on the 200k explanations generated by GPT-4o and Doubao-1.5 for 1 epoch. Conversely, the label-only variant uses only labels (e.g., [[1]]) for training.
After inference on the test set, we decouple matching predictions and explanations through regularization to isolate their respective contributions. We provide a comprehensive evaluation of GLEAM-X through detailed analyses on matching accuracy (Sec.5.2.1) and semantic accuracy (Sec.5.2.2). We further investigate positive/negative sample characteristics (Sec. 5.2.3), explore LLM-as-a-judge and human test beyond Sentence-BERT (Sec. 5.2.4), and discuss potential biases in data curation (Sec. A.7).
@@ 5.2 Evaluation Results on GLEAM-X
Table 6: Positive and negative sample accuracy on GLEAM-X. The Pos-Neg columns report the difference between positive and negative samples. Vanilla Qwen2.5-VL-3B-Instruct shows a strong bias toward predicting non-matching pairs, while model fine-tuning (+GPT-4o, +Doubao-1.5) mitigates this bias.
Model Lang Pos Acc Neg Acc Pos-Neg
GPT-4o 70.83 91.67 -20.84
Doubao-1.5 66.87 88.10 -21.23
Qwen2.5-VL-3B 23.81 90.87 -67.06
     +label only 93.85 91.67 2.18
     +GPT-4o 93.25 87.50 5.75
     +Doubao-1.5 EN 90.08 89.48 0.60
Model Lang Pos Acc Neg Acc Pos-Neg
GPT-4o 73.02 94.25 -21.23
Doubao-1.5 76.19 82.14 -5.95
Qwen2.5-VL-3B 1.19 99.80 -98.61
     +label only 93.85 91.67 2.18
     +GPT-4o 90.87 89.48 1.39
     +Doubao-1.5 ZH 84.52 91.07 -6.55
@@ 5.2.1 Matching Accuracy
Tab. 4 presents the matching accuracy results. Commercial models achieve approximately 80% accuracy, with GPT-4o outperforming Doubao-1.5 on average. In contrast, the original Qwen2.5-VL-3B-Instruct model exhibits significantly lower accuracy, particularly in Chinese (50.50%). After fine-tuning on the explanation dataset, all model variants surpass commercial MLLMs. Notably, the label-only supervision model achieves the highest accuracy, likely because its training targets are extremely short (e.g., [[0]] or [[1]]), reducing learning difficulty. By comparison, the explanation-supervised model must learn the harder task of generating full reasoning while making predictions, resulting in relatively lower accuracy. Despite this, it still substantially outperforms the baseline while providing interpretable explanations for its decisions. This reflects a design choice where GLEAM-X prioritizes interpretable geo-localization over maximizing matching accuracy alone.
@@ 5.2.2 Semantic Analysis of Explanations
Tab. 5 reveals the similarity distributional differences across models. Commercial MLLMs concentrate in the high-similarity range (0.8–1.0) for English, but shift toward 0.6–0.8 in Chinese. The original Qwen2.5-VL-3B places a substantial proportion of predictions in the mid-similarity bins (0.6–0.8) for both languages, indicating weaker semantic alignment with ground-truth explanations. Fine-tuning with LLM-explanation supervision shifts the distribution upward, with both +GPT-4o and +Doubao-1.5 variants peaking in the 0.8–1.0 range. By contrast, the label-only supervision model collapses almost entirely into the lowest bin (0.0–0.2), confirming its inability to generate semantically meaningful explanations despite improvements in matching accuracy.
5.2.3 Positive and Negative Pair Samples
To further analyze model behavior on positive and negative samples, we report matching accuracy in Tab. 6. Vanilla MLLMs, including both commercial models and Qwen2.5-VL-3B-Instruct, exhibit a bias toward negative predictions. Notably, Qwen2.5-VL-3B attains very low accuracy on positive samples while maintaining high accuracy on negative samples, suggesting that the original model tends to classify image pairs as non-matching. This negative bias likely arises from an inherent asymmetry of the task: identifying mismatches can be triggered by a single discrepant cue, whereas confirming a match requires consistent evidence across multiple aspects. After fine-tuning, this bias is substantially alleviated. Positive accuracy increases to over 90% in English and over 84% in Chinese, while negative accuracy remains high. These results indicate that fine-tuning effectively improves the model’s ability to verify matches without sacrificing its capability to detect mismatches.
5.2.4 Discussion: More Semantic Analysis
For semantic analysis, while Sentence-BERT offers low validation costs in terms of both price and computational efficiency, it is limited to measuring textual similarity between model outputs and annotations, potentially failing to capture critical reasoning steps grounded in the visual content itself. To address this limitation, we supplement our evaluation with an LLM-as-a-judge approach using the multimodal Gemini 2.5 Flash (Comanici et al., 2025) model, which can assess both textual reasoning and visual matching. Furthermore, we conduct a human test to validate the effectiveness of both automated assessment methods. Our experiments demonstrate strong alignment across all three evaluation approaches (detailed in Sec. A.6).